# FORWARD: Dataset of a forwarder operating in rough terrain

### v1

@@ Abstract
We present FORWARD, a high-resolution multimodal dataset of a cut-to-length forwarder operating in rough terrain on two harvest sites in the middle part of Sweden. The forwarder is a large Komatsu model equipped with a variety of sensors, including RTK-GNSS, 360-camera, operator vibration sensors, internal CAN-bus signal recording, and multiple IMUs. The data includes event time logs recorded in 5 Hz with e.g., driving speed, fuel consumption, vehicle position with centimeter accuracy, and crane use while the vehicle operates in forest areas laser-scanned with very high-resolution, 1500 points per square meter. Production log files (StanForD standard) with time-stamped machine events, extensive video material, and terrain data in various formats are included as well. About 18 hours of regular wood extraction work during three days is annotated from 360-video material into individual work elements and included in the dataset. We also include scenario specifications of conducted experiments on forest roads and in terrain. Scenarios include repeatedly driving the same routes with and without steel tracks, different load weight, and different target driving speeds. The dataset is intended for developing models and algorithms for trafficability, perception, and autonomous control of forest machines using artificial intelligence, simulation, and experiments on physical testbeds. In part we focus on forwarders traversing terrain, avoiding obstacles, and loading or unloading logs, with consideration for efficiency, fuel consumption, safety, and environmental impact. Other benefits of the open dataset include the ability to explore auto-generation and calibration of forestry machine simulators and automation scenario descriptions using the data recorded in the field. The data, scripts for data exploration and analysis, and some example analyses are long-term publicly available through the Swedish National Data Service.
1 Introduction
@@
The availability of well-annotated data is essential for the development and evaluation of both data-driven and physics-based models essential for automation and high-precision planning. The big advancements in computer vision and deep learning —including image recognition, object detection, and semantic segmentation —is largely a result of the efforts in creating and openly sharing annotated datasets. In autonomous driving, there are several open datasets [1, 2] that support the development of algorithms for localization and mapping, and decision-making [3]. The same data may also be used to create realistic simulators and synthetic generation of data that complements real data and physical testing [4]. In off-road field robotics, terrain traversability is a key functionality for both planning and navigation. Models for predicting the traversability of offroad vehicles and robots from local terrain features is an active field of research and there are several papers with accompanying public datasets [5, 6, 7, 8, 9, 10]. Traversability depends highly on the size, kinematic structure, and driveline of the vehicle or robot [8]. To the best of our knowledge there are no publicly available datasets for articulated heavy off-road vehicles such as a forest machine. This shortage hinders the effective application of artificial intelligence and simulation in research and development towards higher levels of automation and precision forestry, which is needed to achieve sustainability goals. It is the aim of this work to fill this gap in open data.
@@
Creating good datasets is not just a matter of recording large volumes of data, e.g., logging the available data from forest machines in production. With accurate, contextual and consistent annotation, more precise and advanced models can be created from the data, compare [11, 12]. The data needs to have high variability. This may include capturing data from situations that are unusual but safety-critical. Furthermore, a useful dataset needs a well-documented structure as well as instructions and examples for the tools essential for processing the data when it is not in a standardized format.
@@
The FORWARD dataset comprises detailed sensor and log data from a Komatsu cut-to-length forwarder, collected during operations in challenging terrain as well as on gravel roads at two Swedish harvest sites. Sensor instrumentation includes RTK-GNSS for precise positioning, a 360-degree camera for visual documentation, operator seat vibration sensors, CAN-bus signal logging, and multiple IMUs mounted on the machine. The dataset provides time-stamped vehicle positions with centimeter-level accuracy, and machine event logs sampled at 5 Hz, covering variables such as driving speed, fuel consumption, and crane activity. Additionally, high-resolution terrain data from helicopter-borne pre-harvest laser scanning (1500 points/m²), and drone footage of the harvested site is included, along with StanForD production log files containing detailed event records for each work session.
@@
About 18 hours of regular wood extraction work during three days is annotated from 360-video material into individual work elements. We also include controlled experiments, both on forest roads and in terrain. Experiments are structured and presented in a number of scenarios. Scenarios include tests with and without steel tracks, different load weight, and different target driving speeds. The dataset is intended for developing models and algorithms for trafficability, perception, and autonomous control of forest machines using artificial intelligence, simulation, and experiments on physical testbeds. In part we focus on forwarders traversing terrain, avoiding obstacles, and loading or unloading logs, with consideration for efficiency, fuel consumption, safety, and environmental impact. The data, scripts for data exploration and analysis, and some example analyses are publicly available through the Swedish National Data Service [13].
2 Overview of the dataset
@@
An overview of the collected data is provided in Table 1. Details of the data sources, formats, and collection procedure is described in the following sections.
@@
3 Data collection - instrumentation and procedures
Data origins from two field campaigns at two separate sites (Fig 1). Märrviken is the first site with data gathered in October 2023 and very rough terrain with large rocks and boulders. Björsjö is the second site with data gathered in September 2024 and terrain that is less rough but still includes instances of challenging boulders. Terrain data is collected before and during the experiments, using helicopter-borne laser scanning and drone footage-based photogrammetry respectively. The target height of the laser scan with helicopter is 150 meters, resulting in point densities of approximately 1500 points per square meter. A Riegl Vux 120 sensor is used for the laser scanning and the time of scan is July 25-29 2022 and August 19 2023, for Märrviken and Björsjö sites respectively.
During experiments and regular forwarding work at the Björsjö site, we mount a 360°camera on top of the forwarder cabin. The 360°camera records video and audio of the forwarder’s surroundings (Fig 2, left). At the Märrviken site, a few key features were also recorded with a standard mobile phone camera, and converted into 3D models using the online service from Polycam (https://poly.cam).
The forwarder is operated by a professional forwarder operator, who’s seat is equipped with a vibration sensor for parts of the experiments. The vibration sensor, used at the Björsjö site, is a 3D accelerometer with a sampling frequency of 6000 Hz and the vibration dataset is included in this paper. For parts of the experiments at Björsjö, we also mount a number of inertial measurement units (IMU:s) on the forwarder wheels, which record angular velocity at 5 Hz (Fig 2, right). The IMUs are synchronized with the machine data and the 360°video, allowing for detailed analysis of the forwarder’s movement and behavior during the experiments, with specific regard to wheel slip.
4 Data description
4.1 Formats and file structures - time series data
@@
Figure 3: Interactive plot for visualizing and navigating the time series data. The user can zoom in and out, and select preferred time intervals to show experiments, video files, and exact time interval in the terminal (purple box in the figure) for further analysis. Blue lines, orange areas, and green areas indicate experiments, raw 360 video filenames, and lightweight stitched video filenames, respectively.
4.2 Formats and file structures - terrain data
The laser scanning data is provided in LAS format, which is a standard format for storing point cloud data. GeoTIFF files, based on ground classification of the laser scanning data, are provided for each site, with a resolution of between 0.04 and 0.1 m² per pixel. Classified las-files containing only ground points are also provided in LAS format. All three steps from unclassified point cloud, via classified point cloud, to interpolated terrain model are visualized for part of the Björsjö site (Fig 4). Drone images with 60 m flight height are provided as orthomosaic in GeoTIFF format, with a resolution of 2.89 and 1.44 cm² per pixel for Märrviken and Björsjö respectively. The photogrammetric point clouds based on drone footage are provided in LAS format, split into one hectare tiles to reduce individual file size. Each experimental site has its own set of photogrammetric point cloud files accompanied by a metadata CSV-file describing the extent of each tile, easily visualized in geographic information software such as QGIS. Point density for the photogrammetric point clouds are about 5000 and 3000 points/m² for Märrviken and Björsjö respectively. The small 3-D models created from mobile phone footage are provided in OBJ as well as LAZ format, which are standard formats for 3D models.
@@ 4.3 Driving scenario naming convention
Parts of the dataset is based on episodes of recorded machine data during controlled experiments with the forwarder at both the Märrviken and Björsjö sites.
The following naming convention is used to organize the data:
LOCATION_SURFACE_CONFIGURATION-_DESCRIPTION
with
  • positional arguments for location and surface material
    • 3 letter location code per site
    • one-word surface-material description
  • keyword arguments for systematically varied parameters
  • pairs of one-words, separated by underscore: parameter_value
  • principle for ordering: parameters which are more seldom varied are placed first
  • free text descriptions for characterising the test
    • separated by dash
    • multiple words combined with underscore
An example is:
@@ 4.4 Annotation
During the three days of field trails at the Björsjö-site, the forwarder is also operated for 18 hours in regular wood extraction work with the 360°camera mounted on the cabin (Fig 2). This work is annotated into 8 different work elements following established ontology [14], see Table 2. The main objective of the annotation is not to investigate productivity on this specific harvesting site, but to complement the other data with information of work element, e.g., differing between driving with empty or fully loaded bunk (35 vs 55 metric tonnes total machine weight). By analysing annotated work elements together with machine-, position-, and terrain data we also see opportunities to advance automatic work element detection (AWED) in forwarding work. The work elements are defined as follows:
@@
Driving while loading
@@
Driving while unloading
@@
The work element annotations are synchronized to the 360°video material and incorporated in the 5 Hz time series machine data. The active work element can be visualized on the map of the harvesting site using the annotation, exemplified in Fig 14. By utilizing the crane signal data, we can also see when the crane is used, and for how long, during the different work elements. Furthermore, individual load cycles are extracted from the data and used for further visualization and analysis (Fig 15). This is done similarly to the extraction of experiment scenarios, by picking appropriate time intervals from the dataset with the tools provided in the analysis scripts.
@@
5 Analysis tools
Scripts for processing raw data and analysing curated synchronized data are provided in a separate repository available at the Swedish National Data Service (SND) [13]. The scripts are written in Python and use mainly libraries such as NumPy and Matplotlib, along with some specialized libraries for reading and writing specific file formats. Requirements and installation instructions together with usage examples are provided in the repository. The scripts are structured in a modular way, allowing for easy extension and modification. The main functionality is organized in a processing pipeline, where different modules can be combined to perform specific tasks on the data. The pipeline can be constructed and run using a command line interface, optionally in combination with configuration files, with options to specify input and output directories, processing steps, and parameters. The main steps in the processing pipeline are:
Some of the most common data handling and analysis modules are listed in Table 3, a complete list is provided in the repository. Detailed instructions on how to construct and run a number of useful commands are also provided in the repository.
Module Description
TimesToPick Specify start and end times for data extraction
Pick Perform the data extraction based on specified times
Save Save the extracted data in a npz file
PlotAll Plot all loaded variables in the time series data with indexed time on x-axis
Histogram Plot histograms of all loaded variables
6 Results from example analyses
To give more insight into the dataset and exemplify the analysis scripts, we present two example analyses from the high-resolution forwarder dataset and visualization of selected StanForD production data. The first example is an analysis of the fuel consumption during regular work, and the second example is an analysis of the vibration data from both machine and operator seat.
@@ 6.1 Fuel consumption
We test to model the fuel consumption using linear regression with the input variables driving speed, pitch angle, and booleans for full load and tracks on. The resulting model has an $R^{2}$ of 0.62 and a root mean square error (RMSE) of 4.87 l/h. This indicate the possibilities to model fuel consumption, but also that more factors needs to be included when driving in rough terrain (Fig 18). Acknowledging the limitations of this simple model, an interesting observation is that the use of steel tracks seems to affect the fuel consumption more than the load weight, at least in this dataset.

### v2

@@ Abstract
We present FORWARD, a high-resolution multimodal dataset of a cut-to-length forwarder operating in rough terrain on two harvest sites in the middle part of Sweden. The forwarder is a large Komatsu model equipped with vehicle telematics sensors, including global positioning via satellite navigation, movement sensors, accelerometers, and engine sensors. The vehicle was additionally equipped with cameras, operator vibration sensors, and multiple IMUs. The data includes event time logs recorded at 5 Hz of driving speed, fuel consumption, vehicle position with centimeter accuracy, and crane use while the vehicle operates in forest areas, aerially laser-scanned with a resolution of around 1500 points per square meter. Production log files (StanForD standard) with time-stamped machine events, extensive video material, and terrain data in various formats are included as well. About 18 hours of regular wood extraction work during three days is annotated from 360-video material into individual work elements and included in the dataset. We also include scenario specifications of conducted experiments on forest roads and in terrain. Scenarios include repeatedly driving the same routes with and without steel tracks, different load weights, and different target driving speeds. The dataset is intended for developing models and algorithms for trafficability, perception, and autonomous control of forest machines using artificial intelligence, simulation, and experiments on physical testbeds. In part, we focus on forwarders traversing terrain, avoiding or handling obstacles, and loading or unloading logs, with consideration for efficiency, fuel consumption, safety, and environmental impact. Other benefits of the open dataset include the ability to explore auto-generation and calibration of forestry machine simulators and automation scenario descriptions using the data recorded in the field. The data and scripts for data exploration and analysis are made long-term publicly available through the Swedish National Data Service.
Specification Table
Subject Engineering & Materials science
Specific subject area Forest Operations, Automation of Heavy Mobile Equipment, Field Robotics, Machine Learning, Physics-Based Simulation.
Type of data In selection: Vehicle telematics data as time-series (.csv), terrain data (.las, .tif), photos and video data (.360, .mp4, .mov, .jpg), IMU data (.csv), StanForD production data (.hpr, .fpr, .mom), vibration data (.wav), annotations and descriptions in table format (.csv, .xlsx), analysis code (python scripts).
Raw, Filtered, Processed, and with scripts to analyse.
Data collection Data were collected in close connection to regular wood extraction at two harvesting sites. The machine was a large Komatsu forwarder with integrated GNSS-RTK positioning and extra sensors such as wheel-mounted IMU:s, 360-camera, and operator vibration sensor. Sites were LiDAR-scanned (Riegl Vux 120) with helicopter prior to the harvest and photographed with drone post-harvest but pre-extraction.
Data source location The data were collected on the two harvesting sites Märrviken (2023) and Björsjö (2024) in Ånge municipality in the middle of Sweden.
Data accessibility Repository name: Swedish National Data Service (SND), Researchdata.se Data identification number: doi:10.71540/81md-a858 Direct URL to data: https://doi.org/10.71540/81md-a858 Instructions for accessing these data: Review provided README document in the data repository for instructions on download and usage of the data.
Related research article Not yet
1 Value of the data
  • The data provides a comprehensive understanding of the dynamical behavior of a large forest machine traversing rough terrain, passing obstacles, and loading and unloading logs. This is important for developing models and algorithms for perception, planning and autonomous control of sustainable high-precision forestry operations.
  • With centimeter-precision positioning and multiple IMUs, the vehicle motion is captured in detail and can be correlated with the detailed topography of the ground. The same routes are traversed at different speeds and load weights, and with and without steel tracks on the wheels.
  • Researchers can do traversability analysis that goes far beyond today’s standard, which in forestry is limited to the local slope, surface roughness, and bearing capacity. The impact of local surface topography on fuel consumption, whole-body vibrations and effective speed can be studied. This in turn is useful for optimal route planning.
  • The data can be used for validation and calibration of multibody dynamics simulation models which in turn are useful for motion planning and autonomous control of forest machinery.
  • It can be analyzed how experts drive the vehicle and operate the crane in relation to how the logs are distributed over the rough terrain. From this, it is possible to extract automation scenarios and benchmarking data for testing automation solutions and extract statistics on machine utilization in relation to terrain properties.
  • The multimodal data, including high-resolution arial laser scanning, vehicle telematics, video, and production log files have been curated and synchronized. It enables annotation of objects in the terrain (rocks, stumps, logs, vegetation). The data can thus be used for developing refined algorithms for segmenting images or 3D point clouds from forest environments.
  • The data enables detailed analysis of what whole-body vibrations operators are exposed to under different terrain conditions.
2 Background
@@
The availability of well-annotated data is essential for the development and evaluation of both data-driven and physics-based models essential for automation and high-precision planning. The big advancements in computer vision and deep learning—including image recognition, object detection, and semantic segmentation—is largely a result of the efforts in creating and openly sharing annotated datasets. In autonomous driving, there are several open datasets [undef, undefa] that support the development of algorithms for localization and mapping, and decision-making [undefb]. The same data may also be used to create realistic simulators and synthetic generation of data that complements real data and physical testing [undefc]. In off-road field robotics, terrain traversability is a key functionality for both planning and navigation. Models for predicting the traversability of offroad vehicles and robots from local terrain features is an active field of research and there are several papers with accompanying public datasets [undefd, undefe, undeff, undefg, undefh, undefi]. Traversability depends highly on the size, kinematic structure, and driveline of the vehicle or robot [undefg]. To the best of our knowledge there are no publicly available datasets for articulated heavy off-road vehicles such as a forest machine. This shortage hinders the effective application of artificial intelligence and simulation in research and development towards higher levels of automation and precision forestry, which is needed to achieve sustainability goals. It is the aim of this work to fill this gap in open data.
@@
Creating useful datasets is not just a matter of recording large volumes of data, e.g., logging the available data from forest machines in production. With accurate, contextual and consistent annotation, more precise and advanced models can be created from the data, compare [undefj, undefk]. The data needs to have high variability. This may include capturing data from situations that are unusual but safety-critical. Furthermore, a useful dataset needs a well-documented structure as well as instructions and examples for the tools essential for processing the data when it is not in a standardized format.
3 Data description
The FORWARD dataset comprises detailed sensor and log data from a Komatsu cut-to-length forwarder, collected during operations in challenging terrain as well as on gravel roads at two Swedish harvest sites. The complete dataset comprises about 1.1 TB and contain an extensive folder structure and array of different file types, described file-by-file in the accompanying readme-file. In this section, we provide an overview of the data types and their format.
@@ 3 Data description
Sensor instrumentation includes machine-integrated RTK-GNSS for precise positioning, speed, and heading, a 360-degree camera for visual documentation, operator seat vibration sensors, CAN-bus signal logging, and multiple IMUs mounted on the machine. The dataset provides time-stamped vehicle positions with centimeter-level accuracy, and machine event logs sampled at 5 Hz, covering variables such as driving speed, fuel consumption, and crane activity. Additionally, high-resolution terrain data from helicopter-borne pre-harvest laser scanning (1500 points/m2), and drone footage of the harvested site is included, along with StanForD production log files containing detailed event records for each work session.
@@
About 18 hours of regular wood extraction work during three days is annotated from 360-video material into individual work elements. We also include repeated experiments with altered machine configurations, both on forest roads and in terrain. Experiments are structured and presented in a number of scenarios. Scenarios include tests with and without steel tracks, different load weight, and different target driving speeds. The data, scripts for data exploration and analysis, and some example analyses are publicly available through the Swedish National Data Service [undefl].
@@
An overview of the collected data is provided in Table 1. More in-depth information of the data sources and formats are described in the following sections.
@@
3.1 Formats and file structures for time series data
@@
Figure 1: Interactive plot for visualizing and navigating the time series data. The user can zoom in and out, and select preferred time intervals to show experiments, video files, and exact time interval in the terminal (purple box in the figure) for further analysis. Blue lines, orange areas, and green areas indicate experiments, raw 360°-video filenames, and lightweight stitched video filenames, respectively.
3.2 Formats and file structures for terrain data
The laser scanning data is provided in .las format, which is a standard format for storing point cloud data. GeoTIFF (.tif) files, based on ground classification of the laser scanning data, are provided for each site, with a resolution of between 0.04 and 0.1 m2 per pixel. Classified las-files containing only ground points are also provided in .las format. All three steps from unclassified point cloud, via classified point cloud, to interpolated terrain model are visualized for part of the Björsjö site (Fig 2). Drone images with 60 m flight height are provided as orthomosaic in GeoTIFF format, with a resolution of 2.89 and 1.44 cm2 per pixel for Märrviken and Björsjö respectively. The photogrammetric point clouds based on drone footage are provided in .las format, split into one hectare tiles to reduce individual file size. Each experimental site has its own set of photogrammetric point cloud files accompanied by a metadata .csv-file describing the extent of each tile, easily visualized in geographic information software such as QGIS. Point density for the photogrammetric point clouds are about 5000 and 3000 points/m2 for Märrviken and Björsjö respectively. The small 3-D models created from mobile phone footage are provided in .obj and .laz format together with an .mp4 video, all standard formats for 3D models.
@@ 3.2 Formats and file structures for terrain data
3.3 Analysis tools
Scripts for processing raw data and analysing curated synchronized data are provided in a separate repository available at the Swedish National Data Service (SND) [undefl]. The scripts are written in Python and use mainly libraries such as NumPy and Matplotlib, along with some specialized libraries for reading and writing specific file formats. Requirements and installation instructions together with usage examples are provided in the repository. The scripts are structured in a modular way, allowing for easy extension and modification. The main functionality is organized in a processing pipeline, where different modules can be combined to perform specific tasks on the data. The pipeline can be constructed and run using a command line interface, optionally in combination with configuration files, with options to specify input and output directories, processing steps, and parameters. The main steps in the processing pipeline are:
Some of the most common data handling and analysis modules are listed in Table 2, a complete list together with instructions on how to construct and run a number of useful commands is provided in the repository.
Module Description
TimesToPick Specify start and end times for data extraction
Pick Perform the data extraction based on specified times
Save Save the extracted data in a npz file
PlotAll Plot all loaded variables in the time series data with indexed time on horizontal axis
Histogram Plot histograms of all loaded variables
4 Experimental Design, Materials and Methods
Data origins from two field campaigns at two separate sites (Fig 3). Märrviken is the first site with data gathered in October 2023 and very rough terrain with large boulders. Björsjö is the second site with data gathered in September 2024 and terrain that is less rough but still includes instances of challenging boulders. Terrain data is collected before and during the experiments, using helicopter-borne laser scanning and drone footage-based photogrammetry respectively. The target height of the laser scan with helicopter is 150 meters, resulting in point densities of approximately 1500 points per square meter. A Riegl Vux 120 sensor is used for the laser scanning and the time of scan is July 25-29 2022 and August 19 2023, for Märrviken and Björsjö sites respectively.
During experiments and regular forwarding work at the Björsjö site, we mount a 360°-camera on top of the forwarder cabin. The 360°-camera records video and audio of the forwarder’s surroundings (Fig 4, left). At the Märrviken site, a few key features were also recorded with a standard mobile phone camera, and converted into 3D models using the online service from Polycam (https://poly.cam).
The forwarder is operated by a professional forwarder operator, who’s seat is equipped with a vibration sensor for parts of the experiments. The vibration sensor, used at the Björsjö site, is a 3D accelerometer with a sampling frequency of 6000 Hz and the vibration dataset is included in this paper. For parts of the experiments at Björsjö, we also mount a number of inertial measurement units (IMU:s) on the forwarder wheels, which record angular velocity at 5 Hz (Fig 4, right). The IMUs are synchronized with the machine data and the 360°-video, allowing for detailed analysis of the forwarder’s movement and behavior during the experiments, with specific regard to wheel slip.
@@ 4.1 Driving scenario naming convention
Parts of the dataset is based on episodes of recorded machine data during controlled experiments at both the Märrviken and Björsjö sites. The following naming convention is used to organize that data:
LOCATION_SURFACE_CONFIGURATION-DESCRIPTION_1-DESCRIPTION_2
with: positional arguments for location and surface material using three-letter location code per site and one-word surface-material description;keyword arguments for systematically varied parameters; pairs of one-words, separated by underscore (parameter_value); as principle for ordering, parameters which are more seldom varied are placed first; free text descriptions for characterising the test are separated by dash and multiple words combined with underscore. An example is:
@@ 4.2 Annotation
During the three days of field trails at the Björsjö-site, the forwarder is also operated for 18 hours in regular wood extraction work with the 360°-camera mounted on the cabin (Fig 4). This work is annotated into 8 different work elements following established ontology [undefm], see Table 3. The main objective of the annotation is not to investigate productivity on this specific harvesting site, but to complement the other data with information of work element, e.g., differing between driving with empty or fully loaded bunk (35 vs 55 metric tonnes total machine weight). By analysing annotated work elements together with machine-, position-, and terrain data we also see opportunities to advance automatic work element detection (AWED) in forwarding work. The work elements are defined as follows:
[2mm]
@@
Driving while loading:
@@
Driving while unloading:
@@
The work element annotations are synchronized to the 360°-video material and incorporated in the 5 Hz time series machine data. The active work element can be visualized on the map of the harvesting site using the annotation, exemplified in Fig 14. By utilizing the crane signal data, we can also see when the crane is used, and for how long, during the different work elements. Furthermore, individual load cycles are extracted from the data and used for further visualization and analysis (Fig 15). This is done similarly to the extraction of experiment scenarios, by picking appropriate time intervals from the dataset with the tools provided in the analysis scripts.
@@
Applications
To give more insight into the dataset and exemplify applications, we present two concievable use-cases of the dataset and a visualization of selected StanForD production data. The first example is a visualization of the fuel consumption during regular work, and the second example demonstrates the vibration data from the vehicle chassis and the operator seat.
@@ 4.3 Fuel consumption
We demonstrate modeling the fuel consumption using linear regression with the input variables driving speed, pitch angle, and booleans for full load and tracks on. The resulting model has an $R^{2}$ of 0.62 and a root mean square error (RMSE) of 4.87 l/h. This suggests that a fuel consumption model for driving in rough terrain needs to account for more factors than speed, slope, and load weight (Fig 18). An interesting observation is that, in this dataset, the use of steel tracks seems to affect the fuel consumption more than the load weight.